OpenStudy (ashontae19):

-(4h-1) --------- 1+h h=2 I know the answer just i need to know why

OpenStudy (ashontae19):

@chosenmatt

OpenStudy (ashontae19):

do you understand the question i asked?

OpenStudy (anonymous):

no i cant

OpenStudy (ashontae19):

ok hold on

OpenStudy (anonymous):

all right

OpenStudy (ashontae19):

can you see that?

OpenStudy (ashontae19):

@Learner11

OpenStudy (anonymous):

yes. I am trying

OpenStudy (anonymous):

what is your answer?

OpenStudy (ashontae19):

-3

OpenStudy (anonymous):

Sorry , But I got (-7/3) as answer Because , -(4h-1) --------- 1+h then, -[(4*2)-1] -------- 1+2 or, (-8+1)/3 = -7/3

OpenStudy (ashontae19):

but dont you do pemdas ?like -4 *2-1?

OpenStudy (anonymous):

I do . Let me explain

OpenStudy (anonymous):

here , -[(4*2)-1] = -4*2+1 = -8+1 = -7 I was just skip this step before

OpenStudy (ashontae19):

ooo

OpenStudy (ashontae19):

so is all of the explantation?

OpenStudy (anonymous):

XD yes it's all. I have nothing else . Do you not get this?

OpenStudy (ashontae19):

i do but thanks

OpenStudy (ashontae19):

so its -7 pr -7/3

OpenStudy (anonymous):

no prob :D plz give me medal

OpenStudy (anonymous):

thnx .

OpenStudy (ashontae19):

i did

OpenStudy (ashontae19):

is it -7 or -7/3? @Learner11

OpenStudy (anonymous):

no only the -[(4*2)-1] part is -7 as here is also a 3 as divisor So, the answer is -7/3

OpenStudy (anonymous):

see my 5th comment

OpenStudy (ashontae19):

ok thanxs

OpenStudy (anonymous):

:D